Fermi National Laboratory
 

Accelerator Update for June 5, 2009 - June 8, 2009

 

The Integrated Luminosity from6/1/09 to 6/8/09 was 45.62 pb-1.

Friday June 5

The day shift began with vacuum personnel investigating a Tevatron (TeV) cryo system valve problem, with Operations monitoring the Recycler stash (264.4E10), with the Antiproton Source stacking (25.2E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, NuMI, and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store.

Vacuum personnel accessed the TeV and repaired a valve at sector A4.

MTest took beam through the day shift except during problems with a power supply (HP3DS) that lasted from noon to 1:45 PM.

Operations put the TeV into HEP shot setup at 3 PM.

Operations established store 7114 at 6:11 PM with an average initial luminosity (AIL) of 269.2E30.

Operations disabled the MTest beam at 7 PM.

Saturday June 6

The midnight shift began with Operations monitoring store 7114 (125.17E30) and the stash (201.0E10) and the stack (23.2E10), and with NuMI and MiniBooNE taking beam.

Operations established beam to MTest at 4 AM.

Operations terminated store 7114 at 6:14 AM. Operations prepared D0 for a collision hall access.

Operations put the TeV into HEP shot setup at 9:11 AM.

Operations established store 7117 at 11:11 AM with an AIL of 257.04E30.

Sunday June 7

The midnight shift began with Operations monitoring store 7117 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations terminated store 7117 at midnight.

Operations put the TeV into shot setup at 12:10 AM, but the setup was halted due to Recycler sequencer issues.

Operations restarted the TeV shot setup at 2:05 AM.

Operations established store 7119 at 3:33 AM with an AIL of 281.53E30.

Operations established beam to MTest at 4 AM.

Store 7119 aborted at 4:29 AM with QPM indications at sectors A1, A2, and A3. Investigation of the abort lasted through to the end of the midnight shift. The abort was eventually traced to a TeV sector A3 power supply that had a pinhole LCW leak.

EE Support personnel had the sector A3 power supply repaired by 2:10 PM, but then a bad CAMAC power supply, a dipole regulator, and a bad control card (C460) slowed the turn on further.

Operations put the TeV into HEP shot setup at 7:03 PM.

Operations established store 7122 at 9:15 PM with an AIL of 302.26E30.

Monday June 8

The midnight shift began with Operations monitoring store 7122 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established beam to MTest at 4 AM.

At 4:36 AM, the TeV sector F27 water skid tripped off due to a LCW leak in a Pbar Debuncher power supply (D:H717). Pbar will remain off until the completion of this repair.

The Plans for Monday

The plans for today are to repair D:H717 and then resume stacking, stashing.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/08/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory